2. In industries that process joint products, the costs of the raw materials inputs and the sales values of intermediate and final products are often volatile. Change the data area of your worksheet to match the following:

Chapter 11: Applying Excel
Data
Exhibit 11-7 Santa Maria Wool Cooperative
Cost of wool $216,000
Cost of separation process $40,000
Sales value of intermediate products at split-off point:
Undyed coarse wool $135,000
Undyed fine wool $136,000
Undyed superfine wool $62,000
Costs of further processing (dyeing) intermediate products:
Undyed coarse wool $50,000
Undyed fine wool $60,000
Undyed superfine wool $10,000
Sales value of end products:
Dyed coarse wool $167,000
Dyed fine wool $212,000
Dyed superfine wool $98,000

If your formulas are correct, you should get the correct answers to the following questions.

a. What is the overall profit if all intermediate products are processed into final products?


b. What is the financial advantage (disadvantage) from further processing?


c-1. With these new costs and selling prices, what recommendations would you make concerning the company’s operations?


c-2. If your recommendation in part (c-1) is followed, what would be the company’s overall profit?
Answer #1

a)

Statement of profitability if all the intermediate products are processed into final products
Products Undyed coarse wool Undyed fine wool Undyed superfine wool Total for the company
(a) Sale Value   $ 1,67,000 $ 2,12,000 $   98,000 $ 4,77,000
Joint cost apportioned (Note 1) $ 1,03,784 $ 1,04,552 $   47,664 $ 2,56,000
Further processing cost $     50,000 $     50,000 $   10,000 $ 1,10,000
(b) Total Cost of processing $ 1,53,784 $ 1,54,552 $   57,664 $ 3,66,000
Net Profit (a - b) $     13,216 $     57,448 $   40,336 $ 1,11,000

b)

Further Processing
Products Undyed coarse wool Undyed fine wool Undyed superfine wool Total for the company
Sale Value of end products $ 1,67,000 $     2,12,000 $   98,000 $     4,77,000
Less: sale value at split off points $ 1,35,000 $     1,36,000 $   62,000 $     3,33,000
Incremental Revenue from further processing $     32,000 $        76,000 $   36,000 $     1,44,000
Less: Incremental cost of further processing $     50,000 $        60,000 $   10,000 $     1,20,000
Incremental profit/(loss) $   -18,000 $        16,000 $   26,000 $        24,000

The above calculation shows that on further processing of three products, the overall profit of the company will increase by $24,000, although the company will suffer a loss of $18,000 on further processing of Undyed Coarse Wool

c) The best option for the company is to sell Undyed Coarse wool at the split off point i.e. without further processing the product. The other two products Undyed Fine wool and the Undyed Superfine wool should be further processed to optimize the profitability of the company.

d)

Optimal Plan
Products Undyed coarse wool Undyed fine wool Undyed superfine wool Total for the company
(a) Sale Value (Note 2) $    1,35,000 $     2,12,000 $   98,000 $     4,45,000
Joint cost apportioned $    1,03,784 $     1,04,552 $   47,664 $     2,56,000
Further processing cost $        50,000 $   10,000 $        60,000
(b) Total Cost of processing $    1,03,784 $     1,54,552 $   57,664 $     3,16,000
Net Profit (a - b) $        31,216 $        57,448 $   40,336 $     1,29,000

NOTE 1:

Total Joint cost = cost of wool + cost of separation process = $2,16,000 + $40,000 = $ 2,56,000

In absence of any other information, the joint cost of $2,56,000 has been apportioned to the products in the ratio of their sale value at the split off point

Joint cost apportioned to undyed coarse wool = (135000/333000) x 256000 = $1,03,784

Joint cost apportioned to undyed fine wool = (136000 / 333000) x 256000 = $1,04,552

Joinr cost apportioned to undyed superfine wool = (62000/333000) x 256000 = $47,664

NOTE 2: In the optimum plan, the sale value of undyed coarse wool is takes as $1,35,000, the sale value at the split off point